# Revisiting the Berkeley Admissions data: Statistical Tests for Causal Hypotheses

### v1

@@ Abstract
Reasoning about fairness through correlation-based notions is rife with pitfalls. The 1973 University of California, Berkeley graduate school admissions case from Bickel et al. (1975) is a classic example of one such pitfall, namely Simpson’s paradox. The discrepancy in admission rates among male and female applicants, in the aggregate data over all departments, vanishes when admission rates per department are examined. We reason about the Berkeley graduate school admissions case through a causal lens. In the process, we introduce a statistical test for causal hypothesis testing based on Pearl’s instrumental-variable inequalities (Pearl, 1995). We compare different causal notions of fairness that are based on graphical, counterfactual and interventional queries on the causal model, and develop statistical tests for these notions that use only observational data. We study the logical relations between notions, and show that while notions may not be equivalent, their corresponding statistical tests coincide for the case at hand. We believe that a thorough case-based causal analysis helps develop a more principled understanding of both causal hypothesis testing and fairness.
@@ 1 Introduction
In the fall of $1973$, the Graduate Division of the University of California, Berkeley, made admission decisions for $12763$ applicants to its $101$ departments. The admission rate for $8442$ male applicants was approximately $44.2\%$ and for $4321$ female applicants was approximately $34.6\%$. This disparity prompted Bickel et al. (1975) to investigate whether the Graduate Admissions Office discriminated on the basis of sex. The authors found that despite there being a statistically significant disparity in the aggregate data, when each department was examined, the per-department admission rates did not differ significantly between the sexes, thus making this case an instance of Simpson’s paradox. The resolution was that the “proportion of women applicants tends to be high in departments that are hard to get into and low in those that are easy to get into”. The disparity was therefore attributed to societal biases and the authors concluded that there was “no pattern of discrimination on the part of the admissions committee.”
In the fairness literature, the Berkeley graduate admissions case is a canonical example of Simpson’s paradox, which illustrates the limitations of correlation-based fairness notions such as demographic parity, and therefore motivates the need for causal reasoning of fairness. Pearl (2009, Section 4.5.4) analyzes the Berkeley example and frames the conclusion of Bickel et al. (1975) as discerning the direct effect of sex on admissions outcome by conditioning on the mediator, namely department choice. Most works in the fairness literature that mention the Berkeley example follow this analysis, which is predicated on the assumption that the causal model includes no latent confounders while the causal graph is akin to a simple mediation graph with sex being the treatment, department choice being the mediator and the admissions decision being the outcome. However, in both Pearl (2009) and Pearl and Mackenzie (2018), Pearl notes that merely conditioning on department choice might not always be appropriate. In particular, he cites a fascinating exchange between William Kruskal and Peter Bickel in Fairley and Mosteller (1977) where Kruskal objects to the analysis in Bickel et al. (1975) by pointing out that controlling for department leads to erroneuous conclusions if there is a confounder that affects department choice and admissions outcome. To the best of our knowledge, subsequent works that mention the Berkeley example, do not address the latent confounder issue, including Pearl (2009) where the analysis assumes that the common causes are observed. Further, while there are multiple causal fairness notions proposed in the literature, [^1] Some directly inspired by the Berkeley admissions case, for example the path-dependent counterfactual fairness notion in Kusner et al. (2017, Appendix S4). the issue of statistical testing of these fairness notions has received little attention.
@@
For the Berkeley admissions case, we consider multiple fairness notions based on graphical, counterfactual and interventional queries to the family of causal models defined by our causal modeling assumptions, which allows latent confounding between department choice and admissions outcome. For these notions, we develop new statistical tests. One of our key insights is that the graphical notion of fairness can be tested by using the instrumental-variable (IV) inequalities (Pearl, 1995), thus making our proposed statistical test a new test for the IV inequalities. Conversely, any statistical test for the IV inequalities can be used to test for fairness in settings that are analogous to the Berkeley case. In the process, we also prove a result of independent interest, namely the sharpness of the IV inequalities for the case where the instrument and the effect are binary, and the treatment takes any finite number of values. For the Berkeley example, while our proposed fairness notions correspond to different rungs of the causal hierarchy and are in general not equivalent, we show, rather surprisingly, that the tests are equivalent within the IV setting. Although our results are inspired by the Berkeley case, they can also be applied in other analogous settings, e.g. to investigate sex discrimination in awarding distinctions to PhD students (Bol, 2023).
@@ 1.1 Related Work
The question of fairness in decision-making and predictive systems has received increased attention since the past few decades. See Hutchinson and Mitchell (2019); Barocas et al. (2023) for an excellent historical and technical overview, respectively. While attempts at formalizing fairness lead to correlation-based notions such as fairness through unawareness (Dwork et al., 2012), demographic parity, equality of odds (Hardt et al., 2016) etc., purely observational notions of fairness are at odds with each other Chouldechova (2017); Kleinberg et al. (2017) and are prone to erroneous conclusions. On the other hand, observational notions of fairness are readily translated to statistical tests.
Causal analysis tools such as counterfactuals and interventions provide a framework suitable for fairness. As a result, multiple general fairness notions based on counterfactuals were proposed. Kusner et al. (2017) defined a counterfactual fairness notion that required invariance of the distribution of the decision in a given context, with respect to hypothetical changes in the protected attribute. Nabi and Shpitser (2018) and Zhang et al. (2017) consider path-specific effects. Chiappa (2019) proposes a path-specific counterfactual fairness notion and a related notion appears in the appendix of Kusner et al. (2017). Another separate line of work seeks to explain observed disparity through causal discrimination mechanisms (Zhang and Bareinboim, 2018; Plečko and Bareinboim, 2022).
The Berkeley graduate admissions case makes an appearance in multiple papers to motivate the need for causal fairness notions. Kilbertus et al. (2017); Plečko and Bareinboim (2022); Kusner et al. (2017); Chiappa (2019); Berk et al. (2023) are a few among many works. In addition, the Berkeley example also serves as a motivation to introduce path-specific notions given the assumption that the direct effect of sex on admissions outcome is the only ‘unfair’ path. Also, see Barocas et al. (2023) for a critique of this common assumption. Pearl (2009) considers the Berkeley example at length and illustrates the objection to controlling for the mediator by positing an observed confounder.
Despite the fact that most causal fairness works mention the Berkeley example, to the best of our knowledge, no previous work gives a definitive answer to the question of fairness for the Berkeley dataset under unobserved confounding. Kilbertus et al. (2020) discusses the impact of unmeasured confounding under restrictive parametric assumptions. Zhang and Bareinboim (2018); Plečko and Bareinboim (2022) consider fairness models that allow for specific forms of unobserved confounding. Schröder et al. (2024) build on this by providing sensitivity analysis on fairness of prediction models. However, the kinds of unobserved confounding that they allow affects the sensitive attribute which is different from the kind we allow for in the Berkeley dataset.
@@ 2 Preliminaries
We outline a few definitions that follow the formal setup of Bongers et al. (2021).
@@
A Structural Causal Model (SCM) is a tuple $M={\left({V,W,\mathcal{X},f,P}\right)}$ where a) $V,W$ are disjoint, finite index sets of endogenous and exogenous random variables respectively, b) $\mathcal{X}=\prod_{i\in V\cup W}\mathcal{X}_{i}$ is the domain which is a product of standard measurable spaces $\mathcal{X}_{i}$, c) for every $v\in V$, $f_{v}:\mathcal{X}\mapsto\mathcal{X}_{v}$ is a measurable function, called the structural equations, and d) $P{\left({\mathcal{X}_{W}}\right)}=\bigotimes_{w\in W}P(X_{w})$ is the exogenous distribution which is a product of probability distributions $P(X_{w})$ on $\mathcal{X}_{w}$.
@@
For simplicity of exposition, we restrict attention to acyclic SCMs, i.e. SCMs whose causal graph is acyclic (contains no directed cycle $X\rightarrow\cdots Y\rightarrow X$).
@@
Let $M={\left({V,W,\mathcal{X},f,P}\right)}$ be an acyclic SCM, $C\subseteq V$ and $x_{C}\in\mathcal{X}_{C}$. A random variable $X^{\text{do}(x_{C})}:\mathcal{X}_{W}\mapsto\mathcal{X}_{V\backslash C}\times\mathcal{X}_{W}$ is called a potential outcome of $M$ for $x_{C}$ if a) its $W$-component has the exogenous distribution specified by $M$, i.e., $X^{\text{do}(x_{C})}_{W}\sim P{\left({\mathcal{X}_{W}}\right)}$, and b) it satisfies
@@
If $\,C=\emptyset$, we write $X$ instead of $X^{\text{do}(x_{\emptyset})}$. For any $n$, if $C_{1},C_{2},\cdots C_{n}\subseteq V$, then for any $x_{C_{1}}\in{\cal X}_{C_{1}},x_{C_{2}}\in{\cal X}_{C_{2}},\cdots x_{C_{n}}\in{\cal X}_{C_{n}}$,
@@
We make an explicit positivity assumption and define $\mathbb{M}_{\text{IV}+}\triangleq\left\{M\in\mathbb{M}_{\text{IV}}:\forall z,P_{M}(Z=z)>0\right\}$. Causal graphs of SCMs in the IV model class rule out the directed edge $Z\rightarrow Y$ and any latent confounding between $Z$ and $Y$ as well.
@@ 3 Berkeley Case: No Latent Confounding
The set of endogenous variables consists of the protected attribute, namely sex of the applicant, $S$, the department they applied to, $D$, and the decision of the admissions committee, $A$. We assume that $S,A$ are binary variables and $D$ is a discrete-valued variable taking finite number of values, where $S=0,1$ corresponds to male, female applicants, respectively, and $A=0,1$ corresponds to reject and accept, respectively. [^3] The assumption of binary sex is purely for mathematical simplicity. Given that, possibly, societal biases nudge applicants to departments at differing rates depending on their sex, we assume that $S$ affects $D$. Since departments are the primary decision-making units and have different admission rates, we also assume that $D$ affects $A$. The question of whether acceptance decisions discriminate against sex centers around the direct causal effect of $S$ on $A$, and therefore we allow such an effect in the model. We assume the absence of any confounding between the variables (in addition to the absence of any selection bias). The structural equations are given by
@@
where $U_{S},U_{D}$ and $U_{A}$ denote independent exogenous random variables. We denote the family of SCMs parameterized by the functions in (3) and the exogenous distribution as $\mathbb{M}_{\text{no-cf}}$. For $M\in\mathbb{M}_{\text{no-cf}}$, the causal graph $G(M)$ is a directed acyclic graph (DAG), a subgraph of the one in Figure 2(a).
@@ 3.1 Fairness Notions
We define a fairness notion to be a certain condition that is required to be satisfied by a causal model to be deemed fair. These conditions can take the form of observational, interventional, counterfactual or graphical queries on the SCMs in the families of causal models defined by modeling assumptions, in our case, $\mathbb{M}_{\text{no-cf}}$. While the criteria for the fairness notions in Section 3 are phrased in terms of queries corresponding to different rungs of the causal ladder, in our case, any condition can only be tested using observational data.
The investigation of Berkeley’s admission data was initiated on the observation that the well-known fairness notion of demographic parity $P_{M}{\left({A=1\mid S=0}\right)}=P_{M}{\left({A=1\mid S=1}\right)}$ did not hold. This fairness notion is based purely on observational data and we have already noted that it falls prey to Simpson’s paradox. We now present another observational notion of fairness that can be interpreted as a conditional version of demographic parity.
@@
$M\in\mathbb{M}_{\text{no-cf}}$ is fair according to the observational notion of fairness if it belongs to the null hypothesis set
@@
$\left.\implies P_{M}{\left({A=1\mid S=s,D=d}\right)}=P_{M}{\left({A=1\mid D=d}\right)}\right\}.$
Bickel et al. (1975) proposed this notion for the Berkeley data. A valid test for this notion is a conditional independence test for $A\perp\!\!\!\perp S\mid D$. Indeed, the analysis of Bickel et al. (1975) shows that the data contain not enough evidence to reject the null hypothesis that this conditional independence holds, and therefore, concludes fairness.
@@
Pearl (2009, Section 4.5.3) discusses the direct effect in the context of the Berkeley admissions example, where he objects to conditioning for the department and instead proposes intervening on department choice, which corresponds to the controlled direct effect (CDE) (Pearl, 2001) of the ‘treatment’, $S$, on the outcome, $A$, for every value of the mediator, i.e., every department choice $d$.
@@
$M\in\mathbb{M}_{\text{no-cf}}$ is fair according to the interventional notion of fairness if it belongs to
$H^{0}_{\text{no-cf-inter}}\triangleq\left\{M\in\mathbb{M}_{\text{no-cf}}:\forall d,s,P_{M}{\left({A=1\mid\text{do}(S=s),\text{do}(D=d)}\right)}=P_{M}{\left({A=1\mid\text{do}(D=d)}\right)}\right\}.$
Recent analyses of the Berkeley example emphasize counterfactual notions of fairness. In Pearl (2009, Section 4.5.4), Pearl and Mackenzie (2018), Pearl considers a counterfactual quantity, namely the natural direct effect (NDE) (Robins and Greenland, 1992; Pearl, 2001) by motivating a hypothetical experiment where “all female candidates retain their department preferences but change their gender [sex] identification (on the application form) from female to male”. Subsequent causal fairness works (Nabi and Shpitser, 2018; Chiappa, 2019) build on this and propose fairness notions based on known path-specific versions of NDE where the ‘direct path’ from $S$ to $A$ is viewed as ‘unfair’ as opposed to the ‘fair’ path $S\rightarrow D\rightarrow A$. For the Berkeley example, the NDE$(s^{\prime}\rightarrow s)$ is given by
@@
for $s\neq s^{\prime}$. Note that by Pearl’s mediation formula (Pearl, 2001), the above is identified (assuming $\forall d,s,P_{M}{\left({D=d,S=s}\right)}>0$) as
$\sum\limits_{d}\left(P_{M}{\left({A=1\mid D=d,S=s}\right)}-P_{M}{\left({A=1\mid D=d,S=s^{\prime}}\right)}\right)P_{M}{\left({D=d\mid S=s}\right)}.$
@@
Other counterfactual notions of fairness include those by Kusner et al. (2017). The authors define a counterfactual fairness notion that implies demographic parity (see Section E.1.1 for a proof) for the Berkeley example; we have already seen that this particular fairness notion falls prey to Simpson’s paradox. In the appendix, however, they define a path-dependent notion of counterfactual fairness. [^4] This notion is specifically motivated by the Berkeley example. In Section E.1.2 we show that, in our setting, testing for the path-dependent counterfactual fairness notion is equivalent to testing for the conditional independence $A\perp\!\!\!\perp S\mid D$. We now propose an alternate counterfactual notion of fairness and later compare testing of the same.
@@
$H^{0}_{\text{no-cf-ctrf}}\triangleq\left\{M\in\mathbb{M}_{\text{no-cf}}:\forall d,s,P_{M}(A^{\text{do}(S=s,D=d)}=A^{\text{do}(D=d)})=1\right\}.$
The alternate hypotheses are given by the complement of the null hypotheses w.r.t. $\mathbb{M}_{\text{no-cf}}$. Given that the notions are defined on different rungs of the causal hierarchy, it is perhaps not surprising that they are nested accordingly. The assumption of no confounding simplifies the relations as we can prove equivalence of a few notions under positivity. The proof is deferred to Section C.1.
@@
If the data contains enough evidence to reject conditional independence, then the data generating mechanism is unfair w.r.t. the observational notion of fairness. On the other hand, if the data does not contain enough evidence to reject conditional independence, then the data generating mechanism is fair w.r.t. the observational notion of fairness. However, this extrapolation of the outcome of the statistical test on the fairness implications does not hold for the interventional, counterfactual and graphical notions. The following example illustrates that for the graphical notion of fairness, an unfaithful causal model, where $A$ is directly affected by $S$, could satisfy conditional independence.
@@
Let $M\in\mathbb{M}_{\text{no-cf}}$ be defined as $U_{S}\sim\text{Ber}(\frac{1}{2}),U_{A}\sim\text{Ber}(\frac{1}{2}),U_{D}\sim\text{Ber}(\varepsilon)$ where $\varepsilon\in\left[0,\frac{1}{2}\right)$ and $S=U_{S},D=S\oplus U_{D},A=S\oplus D\oplus U_{A}$. Here, $A\perp\!\!\!\perp S\mid D$ but $S$ is a parent of $A$, i.e., $M\in H^{0}_{\text{no-cf-obs}}$, but $M\notin H^{0}_{\text{no-cf-graph}}$.

### v2

@@ Abstract
Reasoning about fairness through correlation-based notions is rife with pitfalls. The 1973 University of California, Berkeley graduate school admissions case from Bickel et al. [1975] is a classic example of one such pitfall, namely Simpson’s paradox. The discrepancy in admission rates among male and female applicants, in the aggregate data over all departments, vanishes when admission rates per department are examined. We reason about the Berkeley graduate school admissions case through a causal lens. In the process, we introduce a statistical test for causal hypothesis testing based on Pearl’s instrumental-variable inequalities [Pearl, 1995]. We compare different causal notions of fairness that are based on graphical, counterfactual and interventional queries on the causal model, and develop statistical tests for these notions that use only observational data. We study the logical relations between notions, and show that while notions may not be equivalent, their corresponding statistical tests coincide for the case at hand. We believe that a thorough case-based causal analysis helps develop a more principled understanding of both causal hypothesis testing and fairness.
@@ 1 Introduction
In the fall of $1973$, the Graduate Division of the University of California, Berkeley, made admission decisions for $12763$ applicants to its $101$ departments. The admission rate for $8442$ male applicants was approximately $44.2\%$ and for $4321$ female applicants was approximately $34.6\%$. This disparity prompted Bickel et al. [1975] to investigate whether the Graduate Admissions Office discriminated on the basis of sex. The authors found that despite there being a statistically significant disparity in the aggregate data, when each department was examined, the per-department admission rates did not differ significantly between the sexes, thus making this case an instance of Simpson’s paradox. The resolution was that the “proportion of women applicants tends to be high in departments that are hard to get into and low in those that are easy to get into”. The disparity was therefore attributed to societal biases and the authors concluded that there was “no pattern of discrimination on the part of the admissions committee.”
In the fairness literature, the Berkeley graduate admissions case is a canonical example of Simpson’s paradox, which illustrates the limitations of correlation-based fairness notions such as demographic parity, and therefore motivates the need for causal reasoning of fairness. Pearl [2009, Section 4.5.4] analyzes the Berkeley example and frames the conclusion of Bickel et al. [1975] as discerning the direct effect of sex on admissions outcome by conditioning on the mediator, namely department choice. Most works in the fairness literature that mention the Berkeley example follow this analysis, which is predicated on the assumption that the causal model includes no latent confounders while the causal graph is akin to a simple mediation graph with sex being the treatment, department choice being the mediator and the admissions decision being the outcome. However, in both Pearl [2009] and Pearl and Mackenzie [2018], Pearl notes that merely conditioning on department choice might not always be appropriate. In particular, he cites a fascinating exchange between William Kruskal and Peter Bickel in Fairley and Mosteller [1977] where Kruskal objects to the analysis in Bickel et al. [1975] by pointing out that controlling for department leads to erroneous conclusions if there is a confounder that affects department choice and admissions outcome. To the best of our knowledge, subsequent works that mention the Berkeley example, do not address the latent confounder issue, including Pearl [2009] where the analysis assumes that the common causes are observed. Further, while there are multiple causal fairness notions proposed in the literature, [^1] Some directly inspired by the Berkeley admissions case, for example the path-dependent counterfactual fairness notion in Kusner et al. [2017, Appendix S4]. the issue of statistical testing of these fairness notions has received little attention.
@@
For the Berkeley admissions case, we consider multiple fairness notions based on graphical, counterfactual and interventional queries to the family of causal models defined by our causal modeling assumptions, which allows latent confounding between department choice and admissions outcome. For these notions, we develop new statistical tests. One of our key insights is that the graphical notion of fairness can be tested by using the instrumental-variable (IV) inequalities [Pearl, 1995], thus making our proposed statistical test a new test for the IV inequalities. Conversely, any statistical test for the IV inequalities can be used to test for fairness in settings that are analogous to the Berkeley case. In the process, we also prove a result of independent interest, namely the sharpness of the IV inequalities for the case where the instrument and the effect are binary, and the treatment takes any finite number of values. For the Berkeley example, while our proposed fairness notions correspond to different rungs of the causal hierarchy and are in general not equivalent, we show, rather surprisingly, that the tests are equivalent within the IV setting. Although our results are inspired by the Berkeley case, they can also be applied in other analogous settings, e.g. to investigate sex discrimination in awarding distinctions to PhD students [Bol, 2023].
@@ 1.1 Related Work
The question of fairness in decision-making and predictive systems has received increased attention since the past few decades. See Hutchinson and Mitchell [2019], Barocas et al. [2023] for an excellent historical and technical overview, respectively. While attempts at formalizing fairness lead to correlation-based notions such as fairness through unawareness [Dwork et al., 2012], demographic parity, equality of odds [Hardt et al., 2016] etc., purely observational notions of fairness are at odds with each other Chouldechova [2017], Kleinberg et al. [2017] and are prone to erroneous conclusions. On the other hand, observational notions of fairness are readily translated to statistical tests.
Causal analysis tools such as counterfactuals and interventions provide a framework suitable for fairness. As a result, multiple general fairness notions based on counterfactuals were proposed. Kusner et al. [2017] defined a counterfactual fairness notion that required invariance of the distribution of the decision in a given context, with respect to hypothetical changes in the protected attribute. Nabi and Shpitser [2018] and Zhang et al. [2017] consider path-specific effects. Chiappa [2019] proposes a path-specific counterfactual fairness notion and a related notion appears in the appendix of Kusner et al. [2017]. Another separate line of work seeks to explain observed disparity through causal discrimination mechanisms [Zhang and Bareinboim, 2018, Plečko et al., 2024].
The Berkeley graduate admissions case makes an appearance in multiple papers to motivate the need for causal fairness notions. Kilbertus et al. [2017], Plečko et al. [2024], Kusner et al. [2017], Chiappa [2019], Berk et al. [2023] are a few among many works. In addition, the Berkeley example also serves as a motivation to introduce path-specific notions given the assumption that the direct effect of sex on admissions outcome is the only ‘unfair’ path. Also, see Barocas et al. [2023] for a critique of this common assumption. Pearl [2009] considers the Berkeley example at length and illustrates the objection to controlling for the mediator by positing an observed confounder.
Despite the fact that most causal fairness works mention the Berkeley example, to the best of our knowledge, no previous work gives a definitive answer to the question of fairness for the Berkeley dataset under unobserved confounding. Kilbertus et al. [2020] discusses the impact of unmeasured confounding under restrictive parametric assumptions. Zhang and Bareinboim [2018], Plečko et al. [2024] consider fairness models that allow for specific forms of unobserved confounding. Schröder et al. [2024] build on this by providing sensitivity analysis on fairness of prediction models. However, the kinds of unobserved confounding that they allow affects the sensitive attribute which is different from the kind we allow for in the Berkeley dataset.
@@ 2 Preliminaries
We outline a few definitions that follow the formal setup of Bongers et al. [2021].
@@
A Structural Causal Model (SCM) is a tuple $M={\left({V,W,\mathcal{X},f,P}\right)}$ where a) $V,W$ are disjoint, finite index sets of endogenous and exogenous random variables respectively, b) $\mathcal{X}=\prod_{i\in V\cup W}\mathcal{X}_{i}$ is the domain which is a product of standard measurable spaces $\mathcal{X}_{i}$, c) for every $v\in V$, $f_{v}:\mathcal{X}\mapsto\mathcal{X}_{v}$ is a measurable function, and $f=(f_{v})_{v\in V}$ is called the causal mechanism; the equations $X_{v}=f_{v}(X)$ for every $v\in V$ are called the structural equations, and d) $P{\left({\mathcal{X}_{W}}\right)}=\bigotimes_{w\in W}P(X_{w})$ is the exogenous distribution which is a product of probability distributions $P(X_{w})$ on $\mathcal{X}_{w}$.
@@
For simplicity of exposition, we restrict attention to acyclic SCMs, i.e. SCMs whose causal graph is acyclic (contains no directed cycle $X\rightarrow\cdots\rightarrow Y\rightarrow X$).
@@
Let $M={\left({V,W,\mathcal{X},f,P}\right)}$ be an acyclic SCM, $C\subseteq V$ and $x_{C}\in\mathcal{X}_{C}$. A random variable $X^{\text{do}(x_{C})}$ taking values in $\mathcal{X}_{V}\times\mathcal{X}_{W}$ is called a potential outcome of $M$ under intervention $x_{C}$ if a) its $W$-component has the exogenous distribution specified by $M$, i.e., $X^{\text{do}(x_{C})}_{W}\sim P{\left({\mathcal{X}_{W}}\right)}$, and b) it satisfies $X^{\text{do}(x_{C})}_{C}=x_{C}\ \mathrm{a.s.}$ and
@@
If $\,C=\emptyset$, we write $X$ instead of $X^{\text{do}(x_{\emptyset})}$. When dealing with multiple potential outcomes, we assume that they share the same $W$-component: for any $n$, for $C_{1},C_{2},\cdots,C_{n}\subseteq V$ and for $x_{C_{1}}\in{\cal X}_{C_{1}},x_{C_{2}}\in{\cal X}_{C_{2}},\cdots,x_{C_{n}}\in{\cal X}_{C_{n}}$,
@@
We make an explicit positivity assumption and define $\mathbb{M}_{\text{IV}+}\triangleq\left\{M\in\mathbb{M}_{\text{IV}}:\forall z,P_{M}(Z=z)>0\right\}$. Causal graphs of SCMs in the IV model class rule out the directed edge $Z\rightarrow Y$ and any latent confounding between $Z$ and $Y$ as well. In Section D.4 we show that our results for the IV model class hold even when confounding is allowed between $Z$ and $X$.
@@ 3 Berkeley Case: No Latent Confounding
The set of endogenous variables consists of the protected attribute, namely sex of the applicant, $S$, the department they applied to, $D$, and the decision of the admissions committee, $A$. We assume that $S,A$ are binary variables and $D$ is a discrete-valued variable taking finite number of values, where $S=0,1$ corresponds to male, female applicants, respectively, and $A=0,1$ corresponds to reject and accept, respectively. [^3] The assumption of binary sex is purely for mathematical simplicity. Given that, possibly, societal biases nudge applicants to departments at differing rates depending on their sex, we assume that $S$ affects $D$. Since departments are the primary decision-making units and have different admission rates, we also assume that $D$ affects $A$. The question of whether acceptance decisions discriminate against sex centers around the direct causal effect of $S$ on $A$ (w.r.t. the three variables $S,D,A$), and therefore we allow such an effect in the model. Indeed, if such a direct causal effect is absent, then the applicant’s sex does not directly affect the admission outcome (only indirectly through department choice). On the other hand, if such a direct causal effect is present, then we would consider this unfair, provided that we know that it is not mediated purely via latent unprotected attributes (e.g., undergraduate department, or undergraduate university chosen by the applicant) that could play a similar ‘fair’ mediating role as $D$. We will—for the time being—assume no latent unprotected attributes exist that mediate an effect of sex on admissions outcome, and come back to this point in Section 6.
In this section, we also assume the absence of any confounding between the variables (in addition to the absence of any selection bias). The structural equations are given by
@@
where $U_{S},U_{D}$ and $U_{A}$ denote independent exogenous random variables. We denote the family of SCMs parameterized by the functions in (3) and the exogenous distribution as $\mathbb{M}_{\text{no-cf}}$. For $M\in\mathbb{M}_{\text{no-cf}}$, the causal graph $G(M)$ is a subgraph of the directed acyclic graph (DAG) in Figure 2(a).
@@ 3.1 Fairness Notions
We define a fairness notion to be a certain condition that is required to be satisfied by a causal model to be deemed fair. These conditions can take the form of constraints expressed in terms of observational, interventional, counterfactual or graphical queries on the SCMs in the families of causal models defined by modeling assumptions, in our case, $\mathbb{M}_{\text{no-cf}}$. While the criteria for the fairness notions in Section 3 are phrased in terms of queries corresponding to different rungs of the causal ladder, in our case, any condition can only be tested using observational data.
The investigation of Berkeley’s admission data was initiated on the observation that the well-known fairness notion of demographic parity $P_{M}{\left({A=1\mid S=0}\right)}=P_{M}{\left({A=1\mid S=1}\right)}$ did not hold. This fairness notion is based purely on observational data and we have already noted that it falls prey to Simpson’s paradox. We now present another observational notion of fairness that can be seen as a conditional version of demographic parity.
@@
$M\in\mathbb{M}_{\text{no-cf}}$ is fair according to the observational notion of fairness if it belongs to
@@
$\!\!\implies\!\!P_{M}{\left({A=1\,|\,S=s,D=d}\right)}=P_{M}{\left({A=1\,|\,D=d}\right)}\}.$
Bickel et al. [1975] proposed this notion for the Berkeley data. It is equivalent to the conditional independence $A\perp\!\!\!\perp S\mid D$. Hence, any valid conditional independence test for $A\perp\!\!\!\perp S\mid D$ provides a statistical test for this notion. Indeed, the analysis of Bickel et al. [1975] shows that the data contain not enough evidence to reject the null hypothesis that this conditional independence holds, and therefore, concludes fairness.
@@
Pearl [2009, Section 4.5.3] discusses the direct effect in the context of the Berkeley admissions example, where he objects to conditioning on the department and instead proposes intervening on department choice, which corresponds to the controlled direct effect (CDE) [Pearl, 2001] of the ‘treatment’, $S$, on the outcome, $A$, for every value of the mediator, i.e., every department choice $d$.
@@
$M\in\mathbb{M}_{\text{no-cf}}$ is fair according to the interventional notion of fairness if it belongs to the null hypothesis set
$H^{0}_{\text{no-cf-inter}}\triangleq\left\{M\in\mathbb{M}_{\text{no-cf}}:\forall d,s\right.$
$\left.P_{M}{\left({A=1\mid\text{do}(S=s),\text{do}(D=d)}\right)}\right.$
$\left.=P_{M}{\left({A=1\mid\text{do}(D=d)}\right)}\right\}.$
Recent analyses of the Berkeley example emphasize counterfactual notions of fairness. In Pearl [2009, Section 4.5.4], Pearl and Mackenzie [2018], Pearl considers a counterfactual quantity, namely the natural direct effect (NDE) [Robins and Greenland, 1992, Pearl, 2001] by motivating a hypothetical experiment where “all female candidates retain their department preferences but change their gender [sex] identification (on the application form) from female to male”. Subsequent causal fairness works [Nabi and Shpitser, 2018, Chiappa, 2019] build on this and propose fairness notions based on known path-specific versions of NDE where the ‘direct path’ from $S$ to $A$ is viewed as ‘unfair’ as opposed to the ‘fair’ path $S\rightarrow D\rightarrow A$. For the Berkeley example, the NDE$(s\rightarrow s^{\prime})$ is given by
@@
for $s\neq s^{\prime}$. Note that by Pearl’s mediation formula [Pearl, 2001], the above is identified (assuming $\forall d,s,P_{M}{\left({D=d,S=s}\right)}>0$) as
$\sum\limits_{d}$ $\left(P_{M}{\left({A=1\mid D=d,S=s^{\prime}}\right)}\right.$
$\left.-P_{M}{\left({A=1\mid D=d,S=s}\right)}\right)P_{M}{\left({D=d\mid S=s}\right)}.$
@@
Other counterfactual notions of fairness include those by Kusner et al. [2017]. The authors define a counterfactual fairness notion that implies demographic parity (see Section E.1.1 for a proof) for the Berkeley example; we have already seen that this particular fairness notion falls prey to Simpson’s paradox. In the appendix, however, they define a path-dependent notion of counterfactual fairness. [^4] This notion is specifically motivated by the Berkeley example. In Section E.1.2 we show that, in our setting, testing for the path-dependent counterfactual fairness notion is equivalent to testing for the conditional independence $A\perp\!\!\!\perp S\mid D$. We now propose an alternate counterfactual notion of fairness and later compare testing of the same.
@@
$H^{0}_{\text{no-cf-ctrf}}$$\triangleq\left\{M\in\mathbb{M}_{\text{no-cf}}:\forall d,s,\right.$
$\left.P_{M}(A^{\text{do}(S=s,D=d)}=A^{\text{do}(D=d)})=1\right\}$
The alternative hypotheses are given by the complement of the null hypotheses w.r.t. $\mathbb{M}_{\text{no-cf}}$. We consider the interventional and counterfactual fairness notions as analogues of the absence of direct effect that are framed in terms of rung-$2$ and rung-$3$ notions, namely interventional and counterfactual distributions, respectively. Therefore, the interventional notion of fairness considers comparing interventional distributions that result from intervening on all the variables with intervening on all variables but the protected variable, and the counterfactual fairness notions compare potential outcomes resulting from the same set of interventions. Given that the notions are defined on different rungs of the causal hierarchy, it is perhaps not surprising that they are nested accordingly. The assumption of no confounding simplifies the relations as we can prove equivalence of a few notions under positivity. The proof is deferred to Section C.1.
@@
If the data contains enough evidence to reject conditional independence, then the data generating mechanism is unfair w.r.t. the observational notion of fairness (assuming no latent unprotected mediators). On the other hand, if the data does not contain enough evidence to reject conditional independence, then the data generating mechanism is fair w.r.t. the observational notion of fairness. However, this extrapolation of the outcome of the statistical test on the fairness implications does not hold for the interventional, counterfactual and graphical notions. The following example illustrates that for the graphical notion of fairness, an unfaithful causal model, where $A$ is directly affected by $S$, could satisfy conditional independence.
@@
Let $M\in\mathbb{M}_{\text{no-cf}}$ be defined as $U_{S}\sim\text{Ber}(\delta),U_{A}\sim\text{Ber}(\varepsilon),U_{D}\sim\text{Ber}(\frac{1}{2})$ where $\delta,\varepsilon\in[0,1]$ and $S=U_{S},D=S\oplus U_{D},A=S\oplus D\oplus U_{A}$. Here, $A\perp\!\!\!\perp S\mid D$ but $S$ is a parent of $A$, i.e., $M\in H^{0}_{\text{no-cf-obs}}$, but $M\notin H^{0}_{\text{no-cf-graph}}$.